I do not know how to set up an equation for this word sentence. I was able to come up with an answer that works in my head by I do not know how to write an equation for it on paper.
The question is: A father is four times as old as his daughter is now. In 20 years, he will be only twice as old as his daugher. How old are the father and daughter now.
The answer I came up with was 7 and 28.
Please help.

Answer by ankor@dixie-net.com(22740)

Here's an easy method to these type of problems:
:
Write an equation for each statement:
:
A father is four times as old as his daughter is now.
F = 4D
:
"In 20 years," their age will be:
F+20
and
D+20
:
In 20 years, he will be only twice as old as his daughter.
F + 20 = 2(D+20)
F + 20 = 2D + 40
F = 2D + 40 - 20
F = 2D + 20
:
How old are the father and daughter now.
From the 1st statement: substitute 4D for F in the above equation:
4D = 2D + 20
4D - 2D = 20
2D = 20
D = 20/2
D = 10 yrs old is the daughter now.
Then
4(10) = 40 yrs old is the father now
:
:
Check our solution using the statement:
"In 20 years, he will be only twice as old as his daughter."
(40 + 20) = 2(10 + 20); confirms out solution
:
Did this help you understand these type problems?